OpenStudy (zolock):

Which is a counterexample that disproves the conjecture? After completing several multiplication problems, a student concludes that the product of two binomials is always a trinomial. (Points : 1) A=(x+7)(x-7) B=(x-1)(x+5) C=(x+3)(x-6) D=(x-5)(x-5)

OpenStudy (anonymous):

\[\left( a+b \right)\left( a-b \right)=a^2-b^2\]

OpenStudy (anonymous):

(x+a)(x-a) =x^2 -a^2

OpenStudy (zolock):

so is it A

OpenStudy (zolock):

can u tell me if its A, B, C, or D

OpenStudy (anonymous):

clearly it is A